Hello! I need some help on two algebra equations and I do have some answers in mind which I will mention at the end. Thank you!!! :)
1. -3(x-14)=42
2. 2-x+4=5

Thanks! Please help, need answers ASAP! Thank you so much!
P.S.
For #1 I thought the answer was x=0 because if you subtract 14 from 0 it would equal -14 and -3*(-14)=42.
For #2 I thought the answer was x=1 because 2-1=1 and 1+4=5. (that one was easy, but hard to use undoing with.)
Hope that helped! Thanks again!!!!

Answer by Alan3354(69443)

1. -3(x-14)=42
2. 2-x+4=5

Thanks! Please help, need answers ASAP! Thank you so much!
P.S.
For #1 I thought the answer was x=0 because if you subtract 14 from 0 it would equal -14 and -3*(-14)=42.
Sub 0 for x
-3(0-14) = 42?
42 = 42 so it's right.
---------------
For #2 I thought the answer was x=1 because 2-1=1 and 1+4=5
2. 2-x+4=5
2 - 1 + 4 = 5?
6-1 = 5?
5 = 5 also correct
=========================
Check them like that.